OpenStudy (daltonw):

Can someone give me a hand? XD 5. Solve for x 16x2 − 81 = 0. x = 4, 9 x = −4, −9 x = , x = − x = , x = −

OpenStudy (daltonw):

ooops it didn't paste correctly

OpenStudy (swamijatin2):

Find the two squares: (4x)^2 – (9)^2 = 0 Factor: (4x + 9)(4x – 9) = 0 Set each of the factors equal to 0, and solve: 4x + 9 = 0; 4x – 9 = 0 x = ± 9/4

OpenStudy (swamijatin2):

did u get the answer

OpenStudy (daltonw):

I can't get the right symboles to show my results! by the way this is a quadratic equation!